# Multiple-policy Evaluation via Density Estimation

### v1

@@ Abstract
In this work, we focus on the multiple-policy evaluation problem where we are given a set of $K$ target policies and the goal is to evaluate their performance (the expected total rewards) to an accuracy $\epsilon$ with probability at least $1-\delta$. We propose an algorithm named $\mathrm{CAESAR}$ to address this problem. Our approach is based on computing an approximate optimal offline sampling distribution and using the data sampled from it to perform the simultaneous estimation of the policy values. $\mathrm{CAESAR}$ consists of two phases. In the first one we produce coarse estimates of the vistation distributions of the target policies at a low order sample complexity rate that scales with $\tilde{O}(\frac{1}{\epsilon})$. In the second phase, we approximate the optimal offline sampling distribution and compute the importance weighting ratios for all target policies by minimizing a step-wise quadratic loss function inspired by the objective in DualDICE Nachum et al. (2019). Up to low order and logarithm terms $\mathrm{CAESAR}$ achieves a sample complexity $\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right)$, where $d^{\pi}$ is the visitation distribution of policy $\pi$ and $\mu^{*}$ is the optimal sampling distribution.
@@ 1 Introduction
Policy evaluation is a fundamental problem in Reinforcement Learning (RL) Sutton and Barto (2018) of which the goal is to estimate the expected total rewards of a given policy. This process serves as an integral component in various RL methodologies, such as policy iteration and policy gradient approaches Sutton et al. (1999), wherein the current policy undergoes evaluation followed by potential updates. Policy evaluation is also paramount in scenarios where prior to deploying a trained policy, thorough evaluation is necessary imperative to ensure its safety and efficacy.
@@
Broadly speaking there exist two scenarios where the problem of policy evaluation has been considered, known as online and offline data regimes. In online scenarios a learner is interacting sequentially with the environment and is tasked with using its online deployments to collect helpful data for policy evaluation. The simplest method for online policy evaluation is Monte-Carlo estimation Fonteneau et al. (2013). One can collect multiple trajectories by following the target policy, and use the empirical mean of the rewards as the estimator. These on-policy methods typically require executing the policy we want to estimate which may be unpractical or dangerous in many cases. For example, in the medical treatment scenario, implementing an untrustworthy policy can cause unfortunate consequences Thapa et al. (2005). In these cases, offline policy evaluation may be preferable. In the offline case, the learner has access to a batch of data and is tasked to use this in the best way possible to estimate the value of a target policy. There are many works focus on this field based on different techniques such as importance-sampling, model-based estimation and doubly-robust estimators Yin and Wang (2020); Jiang and Li (2016); Yin et al. (2021); Xie et al. (2019); Li et al. (2015).
@@
Motivated by the applications where people often have multiple policies that they would like to evaluate, e.g. multiple policies trained using different hyperparameters, Dann et al. (2023) considered multiple-policy evaluation which aims to estimate the performance of a set of $K$ target policies instead of a single policy. From the simplest perspective, multiple-policy evaluation does not pose challenges beyond single-policy evaluation since one can always use single-policy evaluation methods by $K$ times to solve the multiple-policy evaluation problem. However, this can be extremely sample-inefficient as it neglects potential similarities among the $K$ target policies. Consequently, its sample complexity invariably escalates linearly as a function of $K$.
@@
Dann et al. (2023) proposed an on-policy algorithm that leverages the possible similarity among target policies based on an idea related to trajectory synthesis Wang et al. (2020). The basic technique is that if more than one policy take the same action at a certain state, then only one sample is needed at that state which can be reused to synthesize trajectories for these policies. Their algorithm achieves an instance-dependent sample complexity which gives much better results when target policies have many overlaps.
@@
In the context of single policy off-policy evaluation, the theoretical guarantees depend on the overlap between the offline data distribution and the visitations of the evaluated policy Xie et al. (2019); Yin and Wang (2020); Duan et al. (2020). These coverage conditions which ensure that the data logging distribution Xie et al. (2022) adequately covers the state space are typically captured by the ratio between the densities corresponding to the offline data distribution and the policy to evaluate, also known as importance ratios.
@@
A single offline dataset can be used to evaluate multiple policies simultaneously. The policy evaluation guarantees will be different for each of the policies in the set depending on how much overlap the offline distribution has with the policy visitation distributions. These observations inform an approach to the multiple policy evaluation problem different from Dann et al. (2023) that can also leverage the policy visitation overlap in a meaningful way. Our algorithm is based on the idea of designing a behavior distribution with enough coverage of the target policy set. Once this distribution is computed, i.i.d. samples from the behavior distribution can be used to estimate the value of the target policies using ideas inspired in the offline policy optimization literature. Our algorithms consist of two phases:
@@
Coarse estimation of the visitation distributions up to constant multiplicative accuracy can be achieved at a cost that scales linearly, instead of quadratically with the inverse of the accuracy parameter (see Section 4.1) and polynomially in parameters such as the size of the state and action spaces, and the logarithm of the policy evaluation set. Estimating the policy visitation distributions up to multiplicative accuracy is enough to find an approximately optimal behavior distribution that minimizes the maximum visitation ratio among all policies to estimate (see Section 4.2). The samples generated from this behavior distribution are used to estimate the target policy values via importance weighting. Since the importance weights are not known to sufficient accuracy, we propose the $\mathrm{IDES}$ or Importance Density Estimation Algorithm (see Algorithm 1) for estimating these distribution ratios by minimizing a series of loss functions inspired by the $\mathrm{DualDICE}$ Nachum et al. (2019) method (see Section 4.3). Combining these steps we arrive at our main algorithm ($\mathrm{CAESAR}$ ) or Coarse and Adaptive Estimation with Approximate Reweighing for Multi-Policy Evaluation (see Algorithm 2) that achieves a high probability finite sample complexity for the problem of multi-policy evaluation.
@@ 2 Related Work
There is a rich family of off-policy estimators for policy evaluation Liu et al. (2018); Jiang and Li (2016); Dai et al. (2020); Feng et al. (2021); Jiang and Huang (2020). But none of them is effective in our setting. Importance-sampling is a simple and popular method for off-policy evaluation but suffers exponential variance in horizon Liu et al. (2018). Marginalized importance-sampling has been proposed to get rid of the exponential variance. However, existing works all focus on function approximations which only produce approximately correct estimators Dai et al. (2020) or are designed for the infinite-horizon case Feng et al. (2021). Doubly robust estimator Jiang and Li (2016); Hanna et al. (2017); Farajtabar et al. (2018) also solves the exponential variance problem, but no finite sample result is available. Our algorithm is based on marginalized importance-sampling and addresses the above limitations in the sense that our algorithm provides non-asymptotic sample complexity results and works for finite-horizon Markov Decision Processes.
@@
Another popular estimator is called model-based estimator which evaluates the policy by estimating the transition function of the environment Dann et al. (2019); Zanette and Brunskill (2019). Yin and Wang (2020) provides a similar sample complexity to our results. However, there are some significant differences between their result and our result. First, our sampling distribution is optimal which is calculated based on the coarse distribution estimator. Second, our sample complexity is non-asymptotic while their result is asymptotic. Third, the true distributions appear in our sample complexity can be replaced by known distribution estimators without inducing additional costs which means we can provide a known sample complexity while their result is always unknown since we do not know the true visitation distributions of target policies.
@@
The work that most aligns with ours is Dann et al. (2023) which proposed an on-policy algorithm based on the idea of trajectory synthesis. The authors propose the first instance-dependent sample complexity analysis of the multiple-policy evaluation problem. Different from their work, our algorithm uses off-policy evaluation based on importance-weighting and achieves a better sample complexity with simpler techniques and analysis.
@@
Our algorithm also uses some techniques modified from other works which we summarize here. DualDICE is a technique for estimating distribution ratios by minimizing some loss functions proposed by Nachum et al. (2019). We build on this idea and make some modifications to meet the need in our setting. Besides, we utilize stochastic gradient descent algorithms and their convergence rate for strongly-convex and smooth functions in the optimization literature Hazan and Kale (2011). Finally, we adopt the Median of Means estimator Minsker (2023) to convert in-expectation results to high-probability results.
3 Preliminaries and Problem Setup
@@
We denote the set $\{1,2,\dots,N\}$ by $[N]$. $\{X_{n}\}_{n=1}^{N}$ represents the set $\{X_{1},X_{2},\dots,X_{N}\}$. $\mathbb{E}_{\pi}$ denotes the expectation over the trajectories produced by following policy $\pi$. $\tilde{O}$ hides constants, logarithmic and lower-order terms. And we use $\mathbb{V}[X]$ to represent the variance of random variable $X$.
@@
We consider episodic tabular Markov Decision Processes (MDPs) defined by a tuple $\{\mathcal{S},\mathcal{A},H,\{P_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H},\nu\}$ where $\mathcal{S}$ and $\mathcal{A}$ represents the state and action space respectively with $S$ the cardinality of the state space $\mathcal{S}$ and $A$ the cardinality of the action space $\mathcal{A}$. $H$ is the horizon which defines the number of steps the agent can take before the end of an episode. $P_{h}(\cdot|s,a)\in\Delta\mathcal{S}$ is the transition function which represents the probability of transitioning to the next state if the agent takes action $a$ at state $s$. And $r_{h}(s,a)$ is the reward function denotes the reward the agent can get if the agent takes action $a$ at state $s$. In this work, we assume that the reward is deterministic and bounded $r_{h}(s,a)\in[0,1]$ which is consistent with prior work Dann et al. (2023). We denote the initial state distribution by $\nu\in\Delta\mathcal{S}$.
[…]
The state visitation distribution $d^{\pi}_{h}(s)$ of a policy $\pi$ represents the probability of reaching state $s$ at step $h$ if the agent starts from a state sampled from the initial state distribution $\nu$ at step $l=1$ and following policy $\pi$ subsequently, i.e. $d^{\pi}_{h}(s)=\mathbb{P}[s_{h}=s|s_{1}\sim\nu,\pi]$. Similarly, the state-action visitation distribution $d^{\pi}_{h}(s,a)$ is defined as $d^{\pi}_{h}(s,a)=d^{\pi}_{h}(s)\pi(a|s)$. Based on the definition of the visitation distribution, the performance of policy $\pi$ can also be expressed as $J(\pi)=V^{\pi}_{1}=\sum_{h=1}^{H}\sum_{s,a}d^{\pi}_{h}(s,a)r_{h}(s,a)$.
@@
In multiple-policy evaluation, we are given a set of known policies $\{\pi^{k}\}_{k=1}^{K}$ and a pair of factors $\{\epsilon,\delta\}$. The objective is to evaluate the performance of these given policies such that with probability at least $1-\delta$, $\forall\pi\in\{\pi^{k}\}_{k=1}^{K}$, $|\hat{V}^{\pi}_{1}-V^{\pi}_{1}|\leq\epsilon$ where $\hat{V}^{\pi}_{1}$ is the performance estimator.
Dann et al. (2023) proposed an algorithm based on the idea of trajectory stitching and achieved sample complexity,
$\tilde{O}\left(\frac{H^{2}}{\epsilon^{2}}\mathbb{E}\left[\sum_{(s,a)\in\mathcal{K}^{1:H}}\frac{1}{d^{max}(s)}\right]+\frac{SH^{2}K}{\epsilon}\right)$ (1)
where $d^{max}=\max_{k\in[K]}d^{\pi^{k}}$ and $\mathcal{K}^{h}(s,a)\subseteq[K]$ keeps track of which target policies visit state-action pair $(s,a)$ at step $h$ in their trajectories. This sample complexity can be extremely bad due to the existence of $\frac{1}{d^{\max}(s)}$. The authors Dann et al. (2023) also mentioned that their sample complexity is suboptimal by giving a concrete case.
Another way to reuse samples for evaluation of different policies is model-based method. Based on the model-based estimator proposed by Yin and Wang (2020), an asymptotic convergence rate can be derived,
$\sqrt{\frac{H}{n}}\cdot\sqrt{\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}}\left[\frac{d^{\pi^{k}}(s_{h},a_{h})}{\mu(s_{h},a_{h})}\right]}+o(\frac{1}{\sqrt{n}})$ (2)
where $\mu$ is the distribution of the behavior policy used to generate the offline dataset. Though, it looks similar to our results, we have claimed in the section of related work that there are significant differences.

### v2

@@ Abstract
We study the multiple-policy evaluation problem where we are given a set of $K$ policies and the goal is to evaluate their performance (expected total reward over a fixed horizon) to an accuracy $\epsilon$ with probability at least $1-\delta$. We propose an algorithm named $\mathrm{CAESAR}$ for this problem. Our approach is based on computing an approximate optimal offline sampling distribution and using the data sampled from it to perform the simultaneous estimation of the policy values. $\mathrm{CAESAR}$ has two phases. In the first we produce coarse estimates of the visitation distributions of the target policies at a low order sample complexity rate that scales with $\tilde{O}(\frac{1}{\epsilon})$. In the second phase, we approximate the optimal offline sampling distribution and compute the importance weighting ratios for all target policies by minimizing a step-wise quadratic loss function inspired by the DualDICE [21] objective. Up to low order and logarithmic terms $\mathrm{CAESAR}$ achieves a sample complexity $\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right)$, where $d^{\pi}$ is the visitation distribution of policy $\pi$, $\mu^{*}$ is the optimal sampling distribution, and $H$ is the horizon.
@@ 1 Introduction
It is rumored that sometime before the fateful day of March 15, 44 BC, a soothsayer warned Caesar of the impending dangers awaiting him at the Senate house. "Beware the Ides of March," the soothsayer is said to have uttered—a warning that Caesar famously ignored. Immortalized in Shakespeare’s play "Julius Caesar," this cautionary tale reminds us of the dangers of misjudgment and the unforeseen consequences of strategic decisions. In Reinforcement Learning (RL), the stakes, while not mortal, are still significant as agents navigate complex environments to optimize their actions.
@@
This paper delves into the problem of policy evaluation, a fundamental problem in RL [23] of which the goal is to estimate the expected total rewards of a given policy. This process serves as an integral component in various RL methodologies, such as policy iteration and policy gradient approaches [24], wherein the current policy undergoes evaluation followed by potential updates. Policy evaluation is also paramount in scenarios where prior to deploying a trained policy, thorough evaluation is imperative to ensure its safety and efficacy.
@@
Broadly speaking there exist two scenarios where the problem of policy evaluation has been considered, known as online and offline data regimes. In online scenarios, a learner is interacting sequentially with the environment and is tasked with using its online deployments to collect helpful data for policy evaluation. The simplest method for online policy evaluation is Monte-Carlo estimation [11]. One can collect multiple trajectories by following the target policy, and use the empirical mean of the rewards as the estimator. These on-policy methods typically require executing the policy we want to estimate which may be unpractical or dangerous in many cases. For example, in a medical treatment scenario, implementing an untrustworthy policy can cause unfortunate consequences [25]. In these cases, offline policy evaluation may be preferable. In the offline case, the learner has access to a batch of data and is tasked with using it in the best way possible to estimate the value of a target policy. Many works focus on this problem leveraging various techniques, such as importance-sampling, model-based estimation, and doubly-robust estimators [29, 16, 30, 27, 18].
@@
Motivated by applications where one has multiple policies to evaluate, e.g., multiple policies trained using different hyperparameters, Dann et al. [5] considered multiple-policy evaluation which aims to estimate the performance of a set of $K$ target policies instead of a single one. At first glance, multiple-policy evaluation does not pose challenges beyond single-policy evaluation since one can always use $K$ instances of single-policy evaluation to evaluate the $K$ policies. However, since the sample complexity of this procedure scales linearly as a function of $K$ this can be extremely sample-inefficient as it neglects potential similarities among the $K$ target policies.
@@
Dann et al. [5] proposed an on-policy algorithm that leverages the similarity among target policies based on an idea related to trajectory synthesis [26]. The basic technique is that if more than one policy take the same action at a certain state, then only one sample is needed at that state which can be reused to synthesize trajectories for these policies. Their algorithm achieves an instance-dependent sample complexity which gives much better results when target policies have many overlaps.
@@
In the context of single off-policy evaluation, the theoretical guarantees depend on the overlap between the offline data distribution and the visitations of the evaluated policy [27, 29, 8]. These coverage conditions, which ensure that the data logging distribution [28] adequately covers the state space, are typically captured by the ratio between the densities corresponding to the offline data distribution and the policy to evaluate, also known as importance ratios.
@@
A single offline dataset can be used to evaluate multiple policies simultaneously. The policy evaluation guarantees will be different for each of the policies in the set, depending on how much overlap the offline distribution has with the policy visitation distributions. These observations inform an approach to the multiple policy evaluation problem different from [5] that can also leverage the policy visitation overlap in a meaningful way. Our algorithm is based on the idea of designing a behavior distribution with enough coverage of the target policy set. Once this distribution is computed, i.i.d. samples from the behavior distribution can be used to estimate the value of the target policies using ideas inspired in the offline policy optimization literature. Our algorithms consist of two phases:
@@
Coarse estimation of the visitation distributions up to constant multiplicative accuracy can be achieved at a cost that scales linearly, instead of quadratically with the inverse of the accuracy parameter (see Section 4.1) and polynomially in parameters such as the size of the state and action spaces, and the logarithm of the cardinality of the policy evaluation set. We propose the $\mathrm{MARCH}$ or Multi-policy Approximation via Ratio-based Coarse Handling algorithm (see Algorithm 3) for coarse estimation of the visitation distributions. Estimating the policy visitation distributions up to multiplicative accuracy is enough to find an approximately optimal behavior distribution that minimizes the maximum visitation ratio among all policies to estimate (see Section 4.2). The samples generated from this behavior distribution are used to estimate the target policy values via importance weighting. Since the importance weights are not known to sufficient accuracy, we propose the $\mathrm{IDES}$ or Importance Density EStimation algorithm (see Algorithm 1) for estimating these distribution ratios by minimizing a series of loss functions inspired by the DualDICE [21] method (see Section 4.3). Combining these steps we arrive at our main algorithm ($\mathrm{CAESAR}$ ) or Coarse and Adaptive EStimation with Approximate Reweighing algorithm (see Algorithm 2) that achieves a high probability finite sample complexity for the problem of multi-policy evaluation.
@@ 2 Related Work
There is a rich family of off-policy estimators for policy evaluation [19, 16, 4, 10, 15]. But none of them is effective in our setting. Importance-sampling is a simple and popular method for off-policy evaluation but suffers exponential variance in the horizon [19]. Marginalized importance-sampling has been proposed to get rid of the exponential variance. However, existing works all focus on function approximations which only produce approximately correct estimators [4] or are designed for the infinite-horizon case [10]. The doubly robust estimator [16, 13, 9] also solves the exponential variance problem, but no finite sample result is available. Our algorithm is based on marginalized importance-sampling and addresses the above limitations in the sense that it provides non-asymptotic sample complexity results and works for finite-horizon Markov Decision Processes (MDPs).
@@
Another popular estimator is called model-based estimator, which evaluates the policy by estimating the transition function of the environment [6, 31]. Yin and Wang [29] provide a similar sample complexity to our results. However, there are some significant differences between their result and ours. First, our sampling distribution; calculated based on the coarse distribution estimator, is optimal. Second, our sample complexity is non-asymptotic while their result is asymptotic. Third, the true distributions appearing in our sample complexity can be replaced by known distribution estimators without inducing additional costs, that is, we can provide a known sample complexity while their result is always unknown since we do not know the true visitation distributions of target policies.
@@
The work that most aligns with ours is [5], which proposed an on-policy algorithm based on the idea of trajectory synthesis. The authors propose the first instance-dependent sample complexity analysis of the multiple-policy evaluation problem. Different from their work, our algorithm uses off-policy evaluation based on importance-weighting and achieves a better sample complexity with simpler techniques and analysis.
In concurrent work, Amortila et al. [2] propose an exploration objective for downstream reward maximization, similar to our goal of computing an optimal sampling distribution. However, our approximate objective, based on coarse estimation is easier to solve, which is a significant contribution while they need layer-by-layer induction. They also introduced a loss function to estimate ratios, similar to how we estimate the importance densities. However, our ratios are defined differently from theirs which require distinct techniques.
@@
Our algorithm also uses some techniques modified from other works which we summarize here. DualDICE is a technique for estimating distribution ratios by minimizing some loss functions proposed by [21]. We build on this idea and make some modifications to meet the need in our setting. Besides, we utilize stochastic gradient descent algorithms and their convergence rate for strongly-convex and smooth functions from the optimization literature [14]. Finally, we adopt the Median of Means estimator [20] to convert in-expectation results to high-probability results.
3 Preliminaries
@@
We denote the set $\{1,2,\dots,N\}$ by $[N]$. $\{X_{n}\}_{n=1}^{N}$ represents the set $\{X_{1},X_{2},\dots,X_{N}\}$. $\mathbb{E}_{\pi}$ denotes the expectation over the trajectories induced by policy $\pi$. $\tilde{O}$ hides constants, logarithmic and lower-order terms. We use $\mathbb{V}[X]$ to represent the variance of random variable $X$. $\Pi_{det}$ is the set of all deterministic policies and $\text{conv}(\mathcal{X})$ represents the convex hull of the set $\mathcal{X}$.
@@
We consider episodic tabular Markov Decision Processes (MDPs) defined by a tuple $\{\mathcal{S},\mathcal{A},H,\{P_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H},\nu\}$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state and action spaces, respectively, with $S$ and $A$ denoting the respective cardinality of these sets. $H$ is the horizon which defines the number of steps the agent can take before the end of an episode. $P_{h}(\cdot|s,a)\in\Delta\mathcal{S}$ is the transition function which represents the probability of transitioning to the next state if the agent takes action $a$ at state $s$. And $r_{h}(s,a)$ is the reward function, accounting for the reward the agent can collect by taking action $a$ at state $s$. In this work, we assume that the reward is deterministic and bounded $r_{h}(s,a)\in[0,1]$, which is consistent with prior work [5]. We denote the initial state distribution by $\nu\in\Delta\mathcal{S}$.
A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a mapping from the state space to the probability distribution space over the action space. $\pi_{h}(a|s)$ denotes the probability of taking action $a$ at state $s$ and step $h$. The value function $V_{h}^{\pi}(s)$ of a policy $\pi$ is the expected total reward the agent can receive by starting from step $h$, state $s$, and following the policy $\pi$, i.e., $V^{\pi}_{h}(s)=\mathbb{E}_{\pi}[\sum_{l=h}^{H}r_{l}|s]$. The performance $J(\pi)$ of a policy $\pi$ is defined as the expected total reward the agent can obtain. By the definition of the value function, we have $J(\pi)=V_{1}^{\pi}(s|s\sim\nu)$. For simplicity, in the following context, we use $V^{\pi}_{1}$ to denote $V^{\pi}_{1}(s|s\sim\nu)$.
The state visitation distribution $d^{\pi}_{h}(s)$ of a policy $\pi$ represents the probability of reaching state $s$ at step $h$ if the agent starts from a state sampled from the initial state distribution $\nu$ at step $l=1$ and follows policy $\pi$ subsequently, i.e., $d^{\pi}_{h}(s)=\mathbb{P}[s_{h}=s|s_{1}\sim\nu,\pi]$. Similarly, the state-action visitation distribution $d^{\pi}_{h}(s,a)$ is defined as $d^{\pi}_{h}(s,a)=d^{\pi}_{h}(s)\pi(a|s)$. Based on the definition of the visitation distribution, the performance of policy $\pi$ can also be expressed as $J(\pi)=V^{\pi}_{1}=\sum_{h=1}^{H}\sum_{s,a}d^{\pi}_{h}(s,a)r_{h}(s,a)$.
@@
In multiple-policy evaluation, we are given a set of known policies $\{\pi^{k}\}_{k=1}^{K}$ and a pair of factors $\{\epsilon,\delta\}$. The objective is to evaluate the performance of these given policies such that with probability at least $1-\delta$, $\forall\pi\in\{\pi^{k}\}_{k=1}^{K}$, $|\hat{V}^{\pi}_{1}-V^{\pi}_{1}|\leq\epsilon$, where $\hat{V}^{\pi}_{1}$ is the performance estimator.
Dann et al. [5] proposed an algorithm based on the idea of trajectory stitching and achieved an instance-dependent sample complexity,
$\tilde{O}\left(\frac{H^{2}}{\epsilon^{2}}\mathbb{E}\left[\sum_{(s,a)\in\mathcal{K}^{1:H}}\frac{1}{d^{max}(s)}\right]+\frac{SH^{2}K}{\epsilon}\right),$ (1)
where $d^{max}(s)=\max_{k\in[K]}d^{\pi^{k}}(s)$ and $\mathcal{K}^{h}\subseteq\mathcal{S}\times\mathcal{A}$ keeps track of which state-action pairs at step $h$ are visited by target policies in their trajectories.
Another way to reuse samples for evaluating different policies is to estimate the model. Based on the model-based estimator proposed by Yin and Wang [29], an asymptotic convergence rate can be derived,
$\sqrt{\frac{H}{n}}\cdot\sqrt{\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}}\left[\frac{d^{\pi^{k}}(s_{h},a_{h})}{\mu(s_{h},a_{h})}\right]}+o\left(\frac{1}{\sqrt{n}}\right),$ (2)
where $\mu$ is the distribution of the offline dataset and $n$ is the number of trajectories in this dataset. Though, it looks similar to our results, we have claimed in the Section 2 that there are significant differences.